OpenStudy (anonymous):

Given: F(x) = 5x - 6 and G(x) = x - 4 G^-1F^-1 =

OpenStudy (anonymous):

G(x)=x-4*1/GF G(x) = x-4/FG G(x)=FGx-4/FG FGx=FG G(x)+4